# Effect of freestream turbulence on the coherent dynamics of a wind turbine wake

### v1

@@ 4 Effect of freestream turbulence on the important frequencies
In order to understand the effect of FST on the coherent motions at different frequencies, the transverse velocity spectra (obtained using FFT) in the tip region of the near wake ($x/D=0.5$, $y/D=0.55$) are first shown in figure 3(a-d). The symbols in the top right corner correspond to the different FST cases shown in figure 1(a). For all the FST cases, the dominant frequency is $3f_{r}$, where $f_{r}$ is the turbine’s rotational frequency. $3f_{r}$ therefore denotes the frequency of the tip vortices for a three-bladed turbine. Apart from $3f_{r}$, $f_{r}$ is also found to be important and is observed in all cases. Biswas and Buxton [2024a] reported multiple harmonics of $f_{r}$ in the tip region in the absence of freestream turbulence. These harmonics up to $6f_{r}$ and even beyond are clearly observed for case 1. The orange band in the low frequency region $0.15\leq St_{D}\leq 0.4$ shows the typical frequency range of wake meandering reported in previous studies [Medici and Alfredsson, 2008, Okulov et al., 2014, Chamorro et al., 2013]. Note from figure 3(a) that without any freestream turbulence, the spectral energy content in this wake meandering frequency band (WMFB) is negligible in the tip region of the near wake. Contrastingly, for all the FST cases (case 2-4) non-negligible spectral peaks are observed in the WMFB even in the tip region. Especially for case 2a, there is a clear peak at $St_{D}\approx 0.27$. This difference indicates that the the physics underpinning wake meandering for case 1 are likely different from the cases with significant levels of freestream turbulence.
The role of the nacelle in the genesis of wake meandering has recently received some attention [Howard et al., 2015, Biswas and Buxton, 2024a]. Howard et al. [2015] argued that the nacelle’s vortex shedding provides an important perturbation that can trigger wake meandering in the absence of freestream turbulence. Biswas and Buxton [2024b] studied the nature of the wake meandering modes without FST. The wake meandering mode was shown to grow radially with downstream distance and it reached the tip region only after $x\approx 3D$. This justifies the absence of spectral peaks in the WMFB observed in figure 3(a). In the presence of freestream turbulence (cases 2-4), perturbations are available everywhere, unlike case 1 where the nacelle’s vortex shedding is the main source of the perturbation [Howard et al., 2015]. Therefore, with FST, wake meandering is likely to start closer to the turbine. As a result, a non-negligible amount of energy content is present in the WMFB for the FST cases. This will be discussed in more detail in the following paragraphs.
@@
For low $T_{i}$ (case 2a), most of the harmonics of $f_{r}$ are observed in the tip region similar to case 1. However, the amplitudes at these frequencies are weaker except $3f_{r}$, which shows a slightly stronger peak (higher energy content). This could be due to the fact that the freestream velocity for case 2a was slightly higher due to the blockage caused by the grid (see figure 22 in appendix 1), therefore bringing $\lambda_{eff}$ closer to the optimum tip speed ratio (yielding maximum $C_{p}$) as can be seen from figure 15b of Biswas and Buxton [2024a]. For case 3, the harmonics such as $2f_{r}$ and $4f_{r}$ are observed, although their amplitudes are significantly diminished. $6f_{r}$ is no longer observed for case 3. Finally, for the highest $T_{i}$ case, no spectral peaks are evident other than at $f_{r}$ and $3f_{r}$ within the field of view considered here ($0.5\leq x/D\leq 5$). These frequencies can however still be present very close to the turbine blades.
Figure 4: Spectra at $x/D=4$ and $y/D=0$ for the cases (a) 1, (b) 2a, (c) 3, and (d) 4b. The transverse velocity spectra stacked in column (i) are obtained using FFT. The power spectra in the columns (ii) - (v) are obtained using Welch’s PSD estimator having window lengths (ii) 4000, (iii) 3000, (iv) 2000 and (v) 1000. The total length of the data was 5455. The windows had 50% overlap in each case.
Figures 4(i,a-d) show the transverse velocity spectra obtained using FFT at $x/D=4$ and $y/D=0$ for the same four cases of increasing turbulence intensity. The spectra obtained using FFT looked noisy and it was not directly evident if the spectra in the WMFB were discrete or continuous. Therefore, the power spectra were re-computed at the same locations using Welch’s PSD estimator to reduce the noise. The length of the overlapping windows ($w$) was varied during the process to obtain a smoother spectrum. For $w=4000$ (column ii in figure 4) and $w=3000$ (column iii in figure 4), discrete peaks are clearly observed in and around the WMFB. A smaller window size oversimplified the spectra (see columns iv - v in figure 4). Therefore, further work concerning spectral analysis around the WMFB was carried out with a window length of 3000 having a $50\%$ overlap. The no grid case showed a distinct peak in the WMFB at $St_{D}\approx 0.2$ (figure 4(a,ii-iii). Another peak was observed at $St_{D}\approx 0.4$. Although the nacelle’s shedding mode had a similar frequency [Biswas and Buxton, 2024a], it is unlikely to be sustained until $x=4D$. This frequency is therefore more likely to be a harmonic of the wake meandering mode. Note that such harmonics of the main shedding frequency are commonly observed in bluff body wakes [Biswas et al., 2022] further bolstering the link between wake meandering and a global vortex shedding mode from the turbine recently reported by Biswas and Buxton [2024a]. A discrete nature of the spectrum in the WMFB is observed even for the cases with FST, at least for the cases with low to medium $T_{i}$. For the higher $T_{i}$ cases however, the wake meandering region looked increasingly broadband likely indicating a fundamental change in the nature of the wake meandering observed. This will be discussed in more detail in the subsequent paragraphs. Interestingly, the amplitudes of the peaks in the WMFB were stronger in the presence of FST compared to the case without FST (see columns ii-iii in figure 4). This is in line with the previous studies who also reported that the amplitude of wake meandering increased in the presence of FST [Espana et al., 2011, Gambuzza and Ganapathisubramani, 2023].
@@
Figure 5: Dominant frequency maps for the cases (a) 1, (b) 2a, (c) 3, and (d) 4b obtained based on power spectra. The power spectra are obtained using Pwelch’s PSD estimation with a window length of 3000.
@@ 4.1 Frequency maps
For a global view of the important frequencies in different parts of the wake, a dominant frequency map is created similarly to that reported in [Biswas and Buxton, 2024a]. These maps indicate the dominant frequency at a particular spatial location obtained from the power spectrum at that location. The window length used in the PSD estimation was 3000. The map of the non-turbulent case for $\lambda=6$ is shown in figure 5(a) and it looks similar to that reported in [Biswas and Buxton, 2024a] obtained through a different set of experiments. In figure 5(a), there is a small region close to the tip near the rotor where the tip vortices ($3f_{r}$) are the strongest. However, they soon interact with each other producing $f_{r}$, which remains the dominant frequency in the wake shear layer even beyond $5D$. The frequencies related to the tip vortices, especially $f_{r}$, remain dominant in a nearly axisymmetric shell type region apart from the region behind the tower [Biswas and Buxton, 2024a]. This shell type region demarcates the inner wake from the outer background fluid. In the inner wake region, multiple low frequencies dominate in large patches and these frequencies fall in the accepted frequency range of wake meandering. Among these, the largest patch has a Strouhal number $\approx 0.22$ and it initiates from near the nacelle. The low frequency zone as a whole expands radially with downstream distance.
For the cases with FST, the high frequency shell breaks down much earlier and a connection between the background turbulence and the inner wake is established in the vicinity of $x/D=2$. For the low $T_{i}$ case (figure 5(b)), although large patches of low frequencies are observed in the wake, the far wake is dominated by a single frequency $St_{D}\approx 0.29$. This frequency is also present close to the rotor in the tip region and corresponds to the frequency previously identified as $St_{D}=0.27$ using FFT in figure 3(b). The slight difference in the frequency is due to the compromised frequency resolution in the Welch’s PSD estimation. As we increase turbulence intensity, many other low frequencies appear in smaller patches, making it look more broadband and turbulent-like in the inner wake region as well as at the wake’s outer edge.
Figure 6: Power spectra in the freestream (without the turbine) and in the wake, $0.5D$ downstream of the turbine for (a) no grid case, (b) case 2a, (c) case 3 and (d) case 4b obtained using Welch’s PSD estimation with a window length of 3000.
@@ 4.2 The nature and amplitude of wake meandering in the presence of FST
To elucidate the nature of wake meandering in the presence of FST, power spectra are shown in figure 6 with and without the turbine. In each figure set, the left sub figure shows the spectrum in the freestream where the turbine would be placed. The right sub figure shows the spectrum in the turbine wake at a distance of $x/D=0.5$ from the turbine itself. The spectra correspond to the same four cases as discussed previously. The energy content in the freestream spectra is the lowest for the no grid case and it increases for the high $T_{i}$ cases as expected. For all the cases in the turbine wake, low frequency peaks can be observed in the central region (near $y=0$) that correspond to frequencies associated with the nacelle’s vortex shedding ($f_{n}$) and wake meandering ($f_{wm}$). For the no grid case, no signature of wake meandering is found in the wake near the tip region. If looked at more closely, some non-negligible energy content can be observed in the tip region for the cases 3 and 4b. Interestingly, the low $T_{i}$ case (case 2a) showed the strongest signature of wake meandering in the tip region. When considering the freestream spectrum for case 2a, a significant energy content can be observed in a similar frequency range and at a similar spatial location (see the green arrow in figure 6(b)). It appears that these frequencies in the freestream are amplified in the turbine wake. The $f_{wm}$ peak in the tip region was nearly 6 times larger than the strongest peak in the same frequency range available within the freestream for case 2a. Further, it seems that the turbine does not equally amplify all frequencies since energy is available over a broader frequency range in the freestream, while in the wake, the frequency band is narrower. A similar amplification of the freestream frequencies was also observed for the case 2b (not shown here for brevity). These observations are in line with those from recent works which suggest that the turbine has a dynamic response [Mao and Sørensen, 2018, Heisel et al., 2018, Gupta and Wan, 2019] implying some frequencies in the freestream can have a greater impact in altering the wake dynamics, or in other words, some frequencies may promote stronger wake meandering. For the medium or high $T_{i}$ cases although a significant energy content is present in the freestream in the WMFB, the amplification of wake meandering is weaker. It is worthwhile to note that FST can also indirectly induce/influence wake meandering through breaking down the tip vortices [Howard et al., 2015]. Assessing the relative importance of FST in promoting wake meandering directly or indirectly through the tip vortices is non trivial and deserves future investigations. Nevertheless, some evidence of interaction between the tip vortices and wake meandering in the presence of FST is presented in the next paragraph.
Figure 7: Spectra for case 2a at $x/D=0.5$ and (a) $y/D=0.55$ and (b) $y/D=0.53$. Subfigures (c) and (d) show the same for the case 2b.
A closer look at figure 3(b), i.e that for case 2a at $x/D=0.5$ and $y/D=0.55$, reveals a new peak at a frequency slightly higher than $3f_{r}$. To understand the origin of this peak, the spectra are reproduced on a linear scale in figure 7(a). This new frequency was found to be numerically equal to $3f_{r}+f_{wm}$, where $f_{wm}=0.27$ is the wake meandering frequency observed towards the outer edge of the wake. Similarly at $x/D=0.5$ and $y/D=0.53$, another new frequency was found which was numerically equal to $3f_{r}-f_{wm}$. A similar observation was also recorded for the case 2b as shown in figures 7(c-d). This is an excellent example of interaction between the tip vortices and wake meandering. More interestingly, the new frequencies are similar to the secondary frequencies observed in previous studies on multiscale flows involving cylinders/prisms of unequal diameters [Baj and Buxton, 2017, Cicolin et al., 2021, Biswas et al., 2022]. The secondary frequencies were numerically equal to the sum and difference of the primary shedding frequencies of the two cylinders and were shown to be produced due to the non-linear triadic interaction between the primary frequencies. Whether the same is true for the new frequencies observed here will be explored in section 6.
Figure 8: Transverse variation of PSD amplitude averaged over WMFB for different FST cases at (a) $x/D=1$, (b) $x/D=3$ and (c) $x/D=5$. Here $N_{f}$ denotes the number frequencies in the WMFB.
Figure 9: PSD amplitudes averaged over WMFB and transverse distance $y$ for different FST cases.
@@
Next we explore how the amplitude of the spectral peaks corresponding to wake meandering vary with different FST levels. In the presence of FST, it was harder to select a single wake meandering frequency that was dominant over the entire wake region for all the cases. Therefore, the power spectrum at a spatial location was averaged over the frequencies in the WMFB to obtain an average strength of wake meandering. The nacelle’s shedding frequency at $\lambda=6$ was previously found to be around $St_{D}\approx 0.42$ [Biswas and Buxton, 2024a]. Therefore, in order to avoid unwanted contribution from the nacelle’s shedding frequency, the averaging was performed over a frequency range $0.15\leq St_{D}\leq 0.37$ (containing $N_{f}=10$ resolved frequencies) only. The average wake meandering profiles hence created are first shown at $x/D=1$ in figure 8(a) for selected cases. In the central part of the wake, i.e. in the proximity of the nacelle, the wake meandering profile for the no grid case shows the strongest peak. The amplitude thereafter decays towards the edge of the wake. This reaffirms the role of the nacelle in initiating wake meandering when the FST level is negligible. For the other cases with FST, the profiles still show a peak in the central region, although diminished, highlighting that the nacelle still has a role in inducing wake meandering in the central wake region. Unlike the no grid case however, the FST cases show significant wake meandering amplitudes towards the edge of the wake due to the amplification of frequencies from the freestream as discussed previously. Furthermore, note that towards the wake edge, the amplitude is the strongest for the low $T_{i}$ case 2a, not for case 4b which had a higher $T_{i}$ and higher $L_{v}$. This again indicates that $T_{i}$ and $L_{v}$ are not the only important parameters, rather the specific nature of the freestream spectrum is important. Further downstream (figures 8b,c), wake meandering for the no grid case gets weaker while for the other two cases it gets stronger.
To compare all the cases simultaneously, the profiles are further averaged over $y$ and the average amplitudes are shown for all the cases considered here in figure 9. Note that nearer to the turbine, for $x/D\leq 2.5$, the average wake meandering amplitudes are comparable for all the cases. Beyond $x/D=2.5$, the differences become more discernible. Wake meandering is the weakest for the no grid case. Further, the amplitude does not increase monotonically with $T_{i}$. At the farthest streamwise location, case 2a shows almost twice stronger wake meandering compared to most of the other cases.

### v2

@@ 3.1 Important frequencies
In order to understand the effect of FST on the coherent motions at different frequencies, we look at the PSDs of transverse fluctuating velocity ($PSD_{v}$). This is first shown in the tip region ($y\approx 0.5D$) close to the rotor ($x/D=0.5$, $y/D=0.55$) in figure 7 for the different cases (in black). Here the PSDs are estimated using Welch’s method with a window size of $N/2$ (where $N=5456$ was the length of the time series) [Welch, 1967]. This yielded a frequency resolution of 0.0244 hz ($St_{U_{\infty},D}=0.0244$). The spectra at the same location without the turbine is also shown in blue for comparison. The Strouhal number, $St_{U_{b},D}$ is calculated based on the rotor averaged bulk velocity ($U_{b}$) and the turbine diameter ($D$). For all the FST cases, the dominant frequency is $3f_{r}$, where $f_{r}$ is the turbine’s rotational frequency ($\approx$1.87 hz). $3f_{r}$ therefore denotes the frequency of the tip vortices for a three-bladed turbine. Apart from $3f_{r}$, $f_{r}$ is also found to be important and is observed in all cases. The harmonics of $f_{r}$ up to $6f_{r}$ and even beyond are clearly observed for case 1. The red band in the low frequency region $0.15\leq St_{U_{b},D}\leq 0.4$ shows the typical frequency range of wake meandering reported in previous studies [Medici and Alfredsson, 2008, Okulov et al., 2014, Chamorro et al., 2013]. Note that for all the cases with a turbulence generating grid, the spectral energy content in this wake meandering frequency band (WMFB) is higher than the no grid case. More interesting observations can be made by comparing the spectra at the same location without the turbine. For the cases with low FST, the turbine seems to add energy in the WMFB (figures 7(a-c)). Whilst, for medium to high FST (figures 7(d-f)), there is an attenuation of energy in the wake in this low frequency region. Especially for case 2a, there is a clear peak in the vicinity of $St_{U_{b},D}$ that is present both in the freestream and in the wake and is amplified in the wake. This clearly shows how the nature of the inflow turbulence can impact the low frequency wake meandering dynamics in the wake. This will be discussed in more detail in the next sections.
@@
For the low $T_{i}$ cases 2a and 2c, most of the harmonics of $f_{r}$ are still observed in the tip region similar to case 1. For the medium $T_{i}$ case 3b, the harmonics such as $2f_{r}$ and $4f_{r}$ are observed, although their amplitudes are significantly diminished. $6f_{r}$ is no longer observed clearly for case 3b. Finally, for the highest $T_{i}$ cases 4b and 4d, no spectral peaks are evident other than at $f_{r}$ and $3f_{r}$ within the field of view considered here ($0.5\leq x/D\leq 5$). The other frequencies can however still be present very close to the turbine blades.
Figure 9: Spectra at $x/D=4$ and $y/D=0$ obtained using Welch’s PSD estimator having window lengths (a) NN, (b) $N/2$ and (c) $N/4$ for the cases 1, 2c, 3b, 4b and 4d (stacked column wise). Here NN corresponds to the total length of the time series. The windows had 50% overlap. The spectra with and without the turbine are shown in black and blue respectively. The shaded region shows the wake meandering frequency band ($0.15\leq St_{U_{b},D}\leq 0.4$).
Figure 7 showed that $f_{r}$ and $3f_{r}$ are the only two tip vortex related frequencies that could be observed for all the FST cases. Therefore, we explore how the amplitude of these two frequencies vary in the streamwise direction for the different cases. The amplitude was measured in terms of the power contained in a frequency peak ($P_{f}(x,y)$) at a particular location. The power was calculated by integrating the PSD over a $\Delta f=0.14$hz $\Delta St_{U_{\infty},D}=0.14$) which was the best estimate of the width of the spectral peaks associated with tip vortex related frequencies ($f_{r}$, $3f_{r}$ etc). The maximum value of $P_{f}(x,y)$ at a particular $x$ was taken and the streamwise variation of the maximum power ($P_{f_{max}}$) is shown in figure 8. As can be expected, $3f_{r}$ or the tip vortices sustain the furthest downstream for the no grid case. As the FST level is increased they decay faster and closer to the turbine. Interestingly, among the low $T_{i}$ cases (group 2), $3f_{r}$ decays the fastest for case 2a which had the lowest $T_{i}$ and the highest $L_{v}$ in group 2. In fact, for the majority of the cases, the decay of $3f_{r}$ can be seen to be more correlated to $L_{v}$ than $T_{i}$. The formation of $f_{r}$ is more complex and energy can be introduced at this frequency due to for instance the blade eccentricities (very close to the turbine) and due to the merging of the tip vortices further downstream [Abraham and Leweke, 2023, Biswas and Buxton, 2024a]. A preferred dependence on either $T_{i}$ or $L_{v}$ could not observed clearly for $f_{r}$. Nevertheless, it decayed closer to the turbine with increasing FST levels.
Figures 9 shows the PSDs at $x/D=4$ and $y/D=0$ for the five cases (excluding case 2a) of increasing turbulence intensity. At this location, the wake meandering frequency can be expected to become important. Note that the tip vortex related frequencies always showed discrete peaks and therefore their detection was largely insensitive to how the spectrum was calculated. In contrast, the detection of the low frequency peaks in the WMFB (shown by the red shaded region in figure 9) was more sensitive to the choice of the method and the associated parameters, due to their less discrete nature. Therefore, to better understand the nature of the spectra in the WMFB, the spectra were computed with different window sizes in the Welch’s method. In figure 9, the rows a, b and c correspond to window lengths of $N$, $N/2$ and $N/4$ respectively (where $N=5456$ is the length of the time series). For the highest window length, the spectra were relatively noisy, whilst the smallest window length oversimplified the spectra. Therefore, all the analysis related to PSD was carried out with the in-between window length of $N/2$ with a $50\%$ overlap. Nevertheless, clear peaks are observed in the WMFB for all the cases, even with the highest FST levels. Furthermore, the amplitude of the peak in the WMFB for all the FST cases were stronger compared to the no grid case (see row (b) in figure 9). This is in line with the previous studies who also reported that the amplitude of wake meandering increased in the presence of FST [Espana et al., 2011, Gambuzza and Ganapathisubramani, 2023]. The clearest peak was however observed for the no grid case at $St_{U_{b},D}\approx 0.2$. For the other cases, this peak is less coherent or more broadband in nature. This is similar to the observation of Bourhis et al. [2025] who noted a broadening of the wake meandering frequency band for high $T_{i}$. Note that for the no grid case, there is another peak observed at $St_{U_{b},D}\approx 0.4$. This peak was likely a signature of the decaying hub vortex. Although some previous studies have found the hub vortex to become incoherent at around this location [Foti et al., 2016], the relatively larger nacelle used in this study likely creates a stronger hub vortex that sustains further downstream (see for instance figure 4(c,d) of [Biswas and Buxton, 2024b]). The freestream spectra calculated at the same location is also shown in blue. Note that for all the cases the turbine has added energy in the WMFB. The energy difference between the freestream and the wake reduces with increasing FST level. This indicates that in the central region of the far wake, wake meandering is still driven primarily by the turbine, not FST.
@@
Figure 10: Dominant frequency maps for the cases (a) 1, (b) 2c, (c) 3b, (d) 4b and (e) 4d obtained based on the power contained in a narrow frequency band. The power spectra are obtained using Welch’s PSD estimation with a window length of $N/2$ ($N=5456$ being the length of the time series).
@@ 3.2 Frequency maps
For a global view of the important frequencies in different parts of the wake, a dominant frequency map is created similarly to that reported in [Biswas and Buxton, 2024a]. These are shown in figure 10 for the five cases of increasing turbulent intensity. Here, we identified the dominant frequency at a spatial location based on the total power contained around that frequency. The power spectra was integrated over a frequency window $\Delta f$ centred around the frequency in question. If $\Delta f$ is too large there would be nonphysical contributions from different frequencies whilst a $\Delta f$ too small would under predict the power in a particular coherent frequency band. $\Delta f=0.14$ hz ($\Delta St_{U_{\infty},D}=0.14$) was chosen which was the best estimate of the width of the spectral peaks associated with tip vortex related frequencies ($f_{r}$, $3f_{r}$ etc). A slight variation of $\Delta f$ didn’t change the final nature of the maps.
The look of the maps also depends on the number/values of the contours chosen. We showed specific contours with two contours separated by a minimum of $\Delta f$, i.e. the frequencies are binned with a bin size of $\Delta f$. A frequency zone is represented just by the mid frequency in the bin with frequency boundaries $f\pm\Delta f/2$. These frequency boundaries are larger compared to the uncertainty caused by the variation in $U_{b}$ for all the frequencies of importance.
The map of the non-turbulent case for $\lambda=6$ is shown in figure 10(a) and it looks similar to that reported in [Biswas and Buxton, 2024a] obtained through a different set of experiments. In figure 10(a), there is a small region close to the tip near the rotor where the tip vortices ($3f_{r}$) are the strongest. However, they soon interact with each other producing $f_{r}$, which remains the dominant frequency in the wake shear layer even beyond $5D$. The frequencies related to the tip vortices, especially $f_{r}$, remain dominant in a nearly axisymmetric shell type region apart from the region behind the tower [Biswas and Buxton, 2024a]. This shell type region demarcates the inner wake from the outer background fluid. In the inner wake region, multiple low frequencies dominate in large patches and these frequencies fall in the accepted frequency range of wake meandering.
For the cases with FST, the high frequency shell breaks down much closer and a connection between the outer turbulent flow and the inner wake is established much earlier. This has consequences in terms of the sustainment of wake meandering in the far wake for these case as we will discuss. For the highest turbulence intensity cases, many new low frequencies appear in smaller patches, making it look more broadband and turbulent-like in the inner wake region as well as at the wake’s outer edge. The interaction with these frequencies likely makes wake meandering more broadband in the far wake.
Figure 11: PSDs in the freestream (without the turbine) and in the wake, $0.5D$ downstream of the turbine for (a) no grid case, (b) case 2a, (c) case 2c, (d) case 3b, (e) case 4b and (f) case 4d, obtained using Welch’s PSD estimation with a window length of $N/2$.
@@ 3.3 The nature and amplitude of wake meandering in the presence of FST
To elucidate the nature of wake meandering in the presence of FST, power spectral densities are shown in figure 11 with and without the turbine. In each figure set, sub figure (ii) shows the spectrum in the turbine wake at a distance of $x/D=0.5$ from the turbine itself. Sub figure (i) shows the spectrum at the same location with the turbine removed. The spectra correspond to the same six cases (1, 2a, 2c, 3b, 4b and 4d) as discussed previously. The energy content in the freestream spectra is the lowest for the no grid case and it increases for the high $T_{i}$ cases as expected. For all the cases in the turbine wake, low frequency peaks can be observed in the central region (near $y=0$) in the range $0.15\lesssim St_{U_{b},D}\lesssim 0.4$ and $0.4\lesssim St_{U_{b},D}\lesssim 0.6$. The former frequency falls in the wake meandering frequency band defined earlier. The latter matches well with the hub vortex or nacelle shedding frequency reported in several previous studies [Iungo et al., 2013, Viola et al., 2014, Howard et al., 2015, Abraham et al., 2019]. The co-existence of the two frequencies in the vicinity of the nacelle is interesting. It is possible that the lower frequency is a subharmonic of the nacelle’s vortex shedding that grows into wake meandering further downstream. This is in line with [Foti et al., 2016] who showed that the meander profiles in the near wake ($x<2D$) followed the hub vortex. Some energy content is also observed in the frequency range $0.6\lesssim St_{U_{b},D}\lesssim 1$ which is likely associated with the altered vortex shedding from the tower [De Cillis et al., 2021, Biswas and Buxton, 2024a]
Interesting observations can be made by looking at the nature of the spectra in the tip region ($y/D\approx 0.5$). Unlike the no grid case, for the other cases, there is a non-negligible energy content in the WMFB in the tip region. Interestingly, the low $T_{i}$ case (case 2a) showed the highest energy content in the WMFB in the tip region. The spectra without the turbine clearly shows some non-negligible energy content in almost the same frequency range that appears to have been amplified in the wake of the turbine. Curiously, there is more energy in the freestream in the WMFB for some other cases such as the 3b, 4b and 4d. However, there is a clear attenuation of the low frequencies in the wake for these cases as observed in previous studies [Chamorro et al., 2012, Jin et al., 2016]. This is most likely due to the fact that for these cases, the freestream spectra in this region is more multiscale in nature (presence of multiple frequencies) in contrast to more of a single frequency for case 2a. Case 2a therefore represented more of an idealised inflow where there is energy localised in a narrow band. An example of such inflow could be the flow behind a hill, where there is a coherent vortex shedding that might interact with turbines. Now, for case 2a, the turbine is placed $5D$ or $26.6M$ ($M$ being the mesh size) behind the regular grid SRG38, where the flow should be well developed. This frequency is therefore most likely introduced by the modulation of the outer edge of the boundary layer developing on the bottom wall of the flume, although the turbine wasn’t strictly placed within the boundary layer (please see figure 2(b)). As was discussed earlier, the grid SRG38 strongly interacted with the boundary layer. For case 2a, the distance of the rotor from the grid was the highest ($5D$) and the height of the boundary layer was also the highest $\approx 0.63D$ for this case (see figure 2(b)). Therefore, this frequency, although not strictly produced by the grid alone, shows interesting physics in this case.
These observations are in line with those from recent works which suggest that the turbine has a dynamic response [Mao and Sørensen, 2018, Heisel et al., 2018, Gupta and Wan, 2019], i.e. the details of the inflow spectra is important in determining the wake dynamics. It is worthwhile to note that FST can also indirectly induce/influence wake meandering through breaking down the tip vortices [Howard et al., 2015]. Assessing the relative importance of FST in promoting wake meandering directly or indirectly through the tip vortices is non trivial and deserves future investigations. Nevertheless, some evidence of interaction between the tip vortices and wake meandering in the presence of FST is presented in the next paragraphs.
Figure 12: Transverse variation of power in WMFB for different FST cases at (a) $x/D=1$, (b) $x/D=3$ and (c) $x/D=5$.
Figure 13: Streamwise variation of average power (averaged in $y$ direction) in WMFB for different FST cases.
Figure 14: Variation of average power in the WMFB (a, b) in the near wake (averaged between $x/D=1$ and $x/D=2$) and (c, d) in the far wake (averaged between $x/D=4$ and $x/D=5$) for different FST cases.
@@
Next we explore how the total power in the WMFB varied for the different FST cases. The nacelle’s shedding frequency at $\lambda=6$ was previously found to be around $St_{D}\approx 0.42$ [Biswas and Buxton, 2024a]. Therefore, in order to avoid unwanted contributions from the nacelle’s shedding frequency, the power spectra was integrated over a frequency range $0.15\leq St_{U_{b},D}\leq 0.37$ (containing $N_{f}\approx 10$ resolved spectral peaks) only. The wake meandering power profiles hence created are first shown at $x/D=1$ in figure 12(a) for selected cases. In the central part of the wake, i.e. in the proximity of the nacelle, the power profile for the no grid case shows the strongest peak. The amplitude thereafter decays towards the edge of the wake. This reaffirms the role of the nacelle in initiating wake meandering when the FST level is negligible. For the other cases with FST, the profiles still show a peak in the central region, although diminished, highlighting that the nacelle is still the key in inducing wake meandering in the central wake region for the FST cases considered here. Unlike the no grid case however, the FST cases show non-negligible wake meandering power towards the edge of the wake. In line with the observation in figure 11(b), case 2a shows the highest power in the tip region in the near wake. For the mid to high FST cases, although figure 12 shows more power compared to the no grid case, the turbine actually attenuated the power already available in the freestream in the WMFB (as was seen in figure 11). Further downstream (figures 12b,c), the power in the central region diminishes for the no grid case, while that for the other cases is enhanced.
To compare all the cases simultaneously, the power profiles are averaged over $y$ and are shown for all the cases considered here in figure 13. Note that nearer to the turbine, for $x/D\leq 2.5$, the average wake meandering powers are more comparable for all the cases. Beyond $x/D=2.5$, the differences become more discernible. In particular, a monotonic decline is observed for the no grid case. Note that for the no grid case, the low frequency region in the inner wake remains shielded by the shell type region dominated by $f_{r}$ (see figure 10(a)) throughout the streamwise extent of the FOV. For this case wake meandering is likely triggered by the hub vortex/nacelle’s shedding and it follows a trend similar to the decay of the hub vortex [Foti et al., 2016]. For the cases with higher FST level, the average powers of wake meandering are similar to the no grid case in the near wake. However, unlike the no grid case, further downstream, wake meandering is sustained for the other cases. This has to do with the earlier breakdown of the tip vortices as depicted in figure 8. The breakdown of the tip vortices likely provides additional impetus to sustain far field wake meandering. Neunaber et al. [2020] divided the wake into four distinct regions. These included the near wake where the centrline $T_{i}$ decayed (the nacelle’s influence was felt) and immediately downstream, a transition region where the tip vortices break down and the shear layer starts to expand. The extents of these locations depended on the FST levels. Following the same terminology, we can say that for the cases with FST here, the transition region ensues much earlier. For the no grid case, the delayed break up of the tip vortices might eventually give a boost to wake meandering further downstream which is not captured in the current FOV. Whether or not wake the meandering amplitude becomes comparable further downstream is an interesting question and deserves further investigation.
To summarise the results, the spanwise averaged power curves in figure 13 are further averaged over $x$ in the intervals $1<x/D<2$ (near wake) and $4<x/D<5$ (far wake) and shown in figure 14. In the near wake, the average powers do not show any clear trend with either $T_{i}$ or $L_{v}$ and are nearly the same for all the cases. This observation confirms that the primary cause of wake meandering here is the turbine itself, not the FST. Further downstream, for all the cases with FST, the average power of wake meandering is discernibly higher than that of the no grid case. This is most likely related to the earlier breakdown of the tip vortices as we will be discussing in more detail in section 5. The results may indicate that for a ‘realistic’ broadband turbulent inflow, the amplitude of wake meandering do not significantly vary (at least for the FST parameters considered here).
Figure 15: Spectra for case 2a at $x/D=0.5$ and (a) $y/D=0.55$ and (b) $y/D=0.52$. Subfigures (c) and (d) show the same for the case 2b.
A closer look at figure 7(b), i.e that for case 2a at $x/D=0.5$ and $y/D=0.55$, reveals a new peak at a frequency slightly higher than $3f_{r}$. To understand the origin of this peak, the spectra are reproduced on a linear scale in figure 15(a). This new frequency was found to be numerically equal to $3f_{r}+f_{wm}$, where $f_{wm}=0.29$ is the wake meandering frequency observed towards the outer edge of the wake. Similarly at $x/D=0.5$ and $y/D=0.52$, another new frequency was found which was numerically equal to $3f_{r}-f_{wm}$. A similar observation, although weaker, was recorded for the case 2b as well as shown in figures 15(c-d). This is an excellent example of interaction between the tip vortices and wake meandering. More interestingly, the new frequencies are similar to the secondary frequencies observed in previous studies on multiscale flows involving cylinders/prisms of unequal diameters [Baj and Buxton, 2017, Cicolin et al., 2021, Biswas et al., 2022]. The secondary frequencies were numerically equal to the sum and difference of the primary shedding frequencies of the two cylinders and were shown to be produced due to the non-linear triadic interaction between the primary frequencies. Whether the same is true for the new frequencies observed here will be explored in section 5.
Figure 16: $1^{st}$, $2^{nd}$ and final iterations of the mode clustering algorithm for the no grid case (a-c), case 2c (d-f), case 3b (g-i), and case 4b (j-l). The modes indicated by the blue arrows were not selected despite being less damped since the associated mode shapes could not be interpreted as physically meaningful.